# How to Understand Whole Software Repository?

### v1

@@ Abstract.
Recently, Large Language Model (LLM) based agents have advanced the significant development of Automatic Software Engineering (ASE). Although verified effectiveness, the designs of the existing methods mainly focus on the local information of codes, e.g., issues, classes, and functions, leading to limitations in capturing the global context and interdependencies within the software system. From the practical experiences of the human SE developers, we argue that an excellent understanding of the whole repository will be the critical path to ASE. However, understanding the whole repository raises various challenges, e.g., the extremely long code input, the noisy code information, the complex dependency relationships, etc. To this end, we develop a novel ASE method named RepoUnderstander by guiding agents to comprehensively understand the whole repositories. Specifically, we first condense the critical information of the whole repository into the repository knowledge graph in a top-to-down mode to decrease the complexity of repository. Subsequently, we empower the agents the ability of understanding whole repository by proposing a Monte Carlo tree search based repository exploration strategy. In addition, to better utilize the repository-level knowledge, we guide the agents to summarize, analyze, and plan. Then, they can manipulate the tools to dynamically acquire information and generate the patches to solve the real-world GitHub issues. Extensive experiments demonstrate the superiority and effectiveness of the proposed RepoUnderstander. It achieved 18.5% relative improvement on the SWE-bench Lite benchmark compared to SWE-agent.
Automatic Software Engineering (ASE), Agents, Large Language Models (LLMs), Fault Localization, Program Repair, Monte Carlo Tree Search (MCTS)
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Automatic programmingccs: Computing methodologies Multi-agent planning
@@ 1. Introduction
Automating Software Engineering (ASE), which aims to automatically accomplish the Software Engineering (SE) tasks, is a promising and challenging research direction. Recent years, in the ASE domain, Large Language Models (LLMs), especially LLM-based agents, have demonstrated their strong general abilities, e.g., the environment awareness ability (Hong et al., 2023; Wang et al., 2024; Kong et al., 2024), planning & reasoning ability (Cognition, 2023; OpenDevin, 2023; Luo et al., 2024; Wang et al., 2024), tool construction (Zhang et al., 2024a) ability, etc.
@@
More recently, an exemplary milestone termed Devin (Cognition, 2023) explores an end-to-end LLM-based agent system for complex real-world SE tasks (i.e., fix real-world Github issues). Specifically, it plans user requirements, utilizes editor, terminal, and search engine tools for independent decision-making and reasoning, and eventually generates code patches to meet the needs. This innovative approach has garnered considerable attention from the AI and SE communities, notably sparking interest in ASE (Zhang et al., 2024b; Yang et al., [n. d.]). For instance, SWE-agent (Yang et al., [n. d.]) strategically designs an Agent Computer Interface (ACI) to empower SE agents in creating & editing code files, navigating repositories, and executing programs. Additionally, AutoCodeRover (Zhang et al., 2024b) extracts abstract syntax trees in programs, iteratively searches for useful information based on requirements, and generates program patches.
@@
Although the pioneers highlighted the road to advanced ASE and achieved promising performance, their designs, focusing on local information, failed to grasp the global context and intricate interdependencies among functions and classes. For example, SWE-agent (Yang et al., [n. d.]) maintains a context window within a code file that allows the agent to scroll up and down. AutoCodeRover (Zhang et al., 2024b) searches functions or classes within the whole repository. Typically, the code comprising a full logic chain for a specific functionality is not arranged sequentially within a single file; rather, it is logically scattered across multiple folders and files. It is difficult to retrieve all relevant code files among maybe thousands of files in a repository, especially starting only from the text in user requirements. This paper argues that a comprehensive understanding of the whole repository becomes the most critical path to ASE.
@@
Undoubtedly, it is challenging to utilize the vast information of an entire repository within LLM. Firstly, a GitHub repository may contain thousands of code files, making it impractical to include them all in the context windows of LLM. Even if it could, an LLM would struggle to accurately capture the code relevant to the objective within such an extensive context. Secondly, since user requirements are often described in natural language, identifying the relevant code within a repository presents a significant challenge. Thirdly, the intrinsic logic of how the code execution is distinctly different from the sequence of the code text in a file. For instance, the location where a bug triggers an error message and the actual place that requires modification may not be in the same file, yet they are certainly logically connected.
To solve this problem, we propose a novel ASE method named RepoUnderstander by guiding the LLM-based agents to comprehensively understand the whole repository and take advantage of the learned the repository-level knowledge. Imaging the human software engineers are solving project-level issues, they will first overview the repository to ensure a full understanding of the functional modules and dependencies that may be involved. Motivated by this practical insight, we first condense the complex repository-level information by constructing the repository knowledge graph in a top-down manner. Concretely, the repository is organized into a hierarchical structure tree, enabling a clear understanding of the code’s context and scope. Besides, to facilitate comprehensive dependency and interaction analysis, the tree structure is further expanded into a reference graph that captures function call relationships.
@@
Subsequently, we propose a Monte Carlo Tree Search (MCTS) based repository exploration strategy to empower the LLM-based agents the ability of collecting and learning repository-level knowledge. Specifically, the agents first collect the critical information regarding to the SE task on the repository knowledge graph by the explore-and-exploit strategy. Then, by simulating multiple trajectories and evaluating their reward score, our method iteratively narrows down the search space and guide the agents to focus on the most relevant areas. In addition, to better utilize the repository-level knowledge, we guide the agents to summarize, analyze, and plan for the repository information regarding to the SE targets. By these designs, it enables the agents to effectively and efficiently collect and learn task-relevant repository-level knowledge, therefore facilitating the precise fault localization and the informed patch generation. Finally, the agents are instructed to manipulate the API tools to dynamically acquire local information, and fix the real-world issues by generating patches. We demonstrate the superiority and effectiveness of RepoUnderstander via extensive experiments and comprehensive analyses. Besides, through carefully analyses during the experiments, we identify an important problem of the SWE-bench dataset, i.e., missing the necessary interface specification for new feature issues. We propose to fix it and achieve more reliable and effective evaluation for our method and the baseline. The main contributions of this paper are summarized as follows.
@@
  • We highlight the whole repository understanding as the crucial path to ASE and propose a novel agent-based method named RepoUnderstander to solve the challenges.
@@
  • Extensive experiments and analyses demonstrate the superiority and effectiveness of RepoUnderstander  [^1] https://github.com/RepoUnderstander/RepoUnderstander.
@@ 2.1. LLM-based Software Engineering Agents
In recent years, Large Language Model (LLM) based AI agents have advanced the development of automatic software engineering. AI agents improve the capabilities of project-level software engineering (SE) tasks through running environment awareness (Hong et al., 2023; Wang et al., 2024; Kong et al., 2024), planning & reasoning (Wang et al., 2024; Cognition, 2023; OpenDevin, 2023; Luo et al., 2024), and tool construction (Zhang et al., 2024a; Lee et al., 2024). Surprisingly, Devin (Cognition, 2023) is a milestone that explores an end-to-end LLM-based agent system to handle complex SE tasks. Concretely, it first plans the requirements of users, then adopts the editor, terminal and search engine tools to make independent decisions and reasoning, and finally generates codes to satisfy the needs of users in an end-to-end manner. Its promising designs and performance swiftly ignited unprecedented attention from the AI community and SE community to Automatic Software Engineering (ASE) (Yang et al., [n. d.]; Zhang et al., 2024b). For example, SWE-agent (Yang et al., [n. d.]) carefully designs an Agent Computer Interface (ACI) to empower the SE agents capabilities of creating & editing code files, navigating repositories, and executing programs. Besides, AutoCodeRover (Zhang et al., 2024b) extracts the abstract syntax trees in programs, then iteratively searches the useful information according to requirements, and eventually generates program patches. Their designs mainly focus on the local information in the repository, e.g., code files, classes, or functions themselves. Although achieving promising performance, from the insights of the human SE developers, the excellent understanding of the whole repository is a critical path to ASE.
@@ 2.2. Evaluation of LLM-based Software Engineering Agents
Benefiting from the strong general capability of LLMs, LLM-based software engineering agents can handle many important SE tasks, e.g., repository navigation (Zhang et al., 2024a; Wang et al., 2024), code generation (Hong et al., 2023; Ding et al., 2024a; Ishibashi and Nishimura, 2024; Tang et al., 2024; Rasheed et al., 2024), debugging (Hong et al., 2023; Yang et al., [n. d.]; Zhang et al., 2024b), program repair (Qin et al., 2024; Zhang et al., 2024b; Yang et al., [n. d.]). The existing methods usually regard code generation as a core ability and mainly conduct evaluations on it. Precisely, the code generation test set (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024; Zheng et al., 2023; Lu et al., 2021) consists of the short problem description, the solution, and the corresponding unit test data. However, with the fast development of LLMs and agents, these datasets are no longer able to comprehensively evaluate their capabilities in the real-world SE tasks. To this end, the repository-level code completion and generation tasks (Liu et al., 2023; Ding et al., 2024b; Du et al., 2024) are presented to evaluate the repository understanding and generation capabilities of LLMs and agents. More recently, SWE team(Jimenez et al., 2024; Yang et al., [n. d.]) develop a unified dataset named SWE-bench to evaluate the capability of the agent system to solve real-world GitHub issues automatically. Specifically, it collects the task instances from real-world GitHub issues from twelve repositories. Consistent with previous evaluation methods, SWE-bench is based on the automatic execution of the unit tests. Differently, the presented test set is challenging and requires the agents to have multiple capabilities, including repository navigation, fault locating, debugging, code generation and program repairing. Besides, SWE-bench Lite (Carlos E. Jimenez, John Yang, Jiayi Geng, 2024) is a subset of SWE-bench, and it has a similar diversity and distribution of repositories as the original version. Due to the smaller test cost and more detailed filtering, SWE-bench Lite is officially recommended as the benchmark of LLM-based SE agents. Therefore, consistent with previous methods (Yang et al., [n. d.]; Zhang et al., 2024b; OpenDevin, 2023), we report our performance on SWE-bench Lite.
@@ 3.1. Overview
We first describe the overall operating process of RepoUnderstander, and introduce the stages in detail in the subsequent parts of this section. Given a workspace, RepoUnderstander can automatically solve real-world GitHub issues. Among them, RepoUnderstander involves three key steps, repository knowledge graph construction stage, MCTS-enhanced repository understanding stage, information utilization & patch generation stage. The overall workflow is shown in Figure 1.
@@
Figure 1. The overview of our proposed RepoUnderstander. Firstly, we construct the repository knowledge graph is constructed to efficiently represent the code and the dependency in the repository. Subsequently, we empower the agents with the ability of repository understanding by designing the Monte Carlo tree search based repository explore strategy. In addition, we guide the agents to summarize, analyze, and plan to better utilize the repository-level knowledge. Then, they can manipulate the tools to dynamically acquire issue-relevant code information and generate the patches to solve the real-world GitHub issues
@@
In Repository Knowledge Graph Construction phase, RepoUnderstander first builds a repository knowledge graph to represent the entire repository and describe the relationships between entities. This is achieved by parsing the software structure and analyzing it in a top-down manner. The repository is first organized into a hierarchical tree that allows a clear understanding of the context and scope of the code. To facilitate comprehensive dependency and interaction analysis, the tree structure is further extended into a reference graph that captures function call relationships.
@@
Due to the large scale and information complexity of the repository knowledge graph, during the MCTS-Enhanced Repository Understanding phase, RepoUnderstander uses the Monte Carlo tree search algorithm to dynamically explore the entire graph. This method focus on discovering key information (i.e., repository functionality and dependency structure) that has a significant impact on issue solving. Through correlation expansion and reference expansion, MCTS simulates multiple trajectories and evaluates their importance, dynamically narrowing the search space and allocating computing resources to the most relevant regions. This targeted navigation enables the model to efficiently access and process important information in the repository, thereby facilitating precise fault localization and informed patch generation.
@@
Inspired by the actual development experience of human programmers, it is necessary to have certain global prior knowledge of the repository before solving specific tasks. Therefore, in Information Utilization & Patch Generation phase, RepoUnderstander first summarizes the important information collected in the repository understanding phase to form an experience of the entire repository. Then, in order for RepoUnderstander to use the global experience to obtain dynamic information during the problem solving process, we follow AutoCodeRover(Zhang et al., 2024b)’s information retrieval method and use API tools to extract information in the repository knowledge graph. This includes specific classes and functions and code snippets, etc., to maintain local dynamic knowledge during the task. After collecting enough context, RepoUnderstander uses global experience to summarize the currently acquired information to locate faults, generate modified code and return patches that try to resolve the issue.

### v2

# Alibaba LingmaAgent: Improving Automated Issue Resolution via Comprehensive Repository Exploration

@@ Abstract.
This paper presents Alibaba LingmaAgent, a novel Automated Software Engineering method designed to comprehensively understand and utilize whole software repositories for issue resolution. Deployed in TONGYI Lingma, an IDE-based coding assistant developed by Alibaba Cloud, LingmaAgent addresses the limitations of existing LLM-based agents that primarily focus on local code information. Our approach introduces a top-down method to condense critical repository information into a knowledge graph, reducing complexity, and employs a Monte Carlo tree search based strategy enabling agents to explore and understand entire repositories. We guide agents to summarize, analyze, and plan using repository-level knowledge, allowing them to dynamically acquire information and generate patches for real-world GitHub issues. In extensive experiments, LingmaAgent demonstrated significant improvements, achieving an 18.5% relative improvement on the SWE-bench Lite benchmark compared to SWE-agent. In production deployment and evaluation at Alibaba Cloud, LingmaAgent automatically resolved 16.9% of in-house issues faced by development engineers, and solved 43.3% of problems after manual intervention. Additionally, we have open-sourced a Python prototype of LingmaAgent for reference by other industrial developers  [^1] https://github.com/RepoUnderstander/RepoUnderstander. In fact, LingmaAgent has been used as a developed reference by many subsequently agents.
Automatic Software Engineering (ASE), Software Engineering Agents, Large Language Models (LLMs), Fault Localization, Automated Program Repair, Monte Carlo Tree Search (MCTS)
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 23–27, 2025; Trondheim, Norwaybooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwayisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Automatic programmingccs: Computing methodologies Multi-agent planning
@@ 1. Introduction
Automated Software Engineering (ASE) explores the automation of complex software development processes and develops innovative tools to improve software lifecycle. Recent years, in the ASE domain, LLM-based agents have demonstrated their strong general abilities, e.g., the environment awareness ability (Hong et al., 2023; Wang et al., 2024a; Kong et al., 2024), planning & reasoning ability (Cognition, 2023; OpenDevin, 2023; Luo et al., 2024; Wang et al., 2024a), tool construction (Zhang et al., 2024a) ability, etc.
@@
More recently, an exemplary milestone termed Devin (Cognition, 2023) explores an end-to-end LLM-based agent system for complex real-world SE tasks (i.e., fix real-world Github issues). It plans user requirements, utilizes editor and terminal tools for independent decision-making and reasoning, and eventually generates code patches to meet the needs. This innovative approach has garnered considerable attention from the AI and SE communities (Zhang et al., 2024b; Yang et al., 2024). For instance, SWE-agent (Yang et al., 2024) strategically designs an Agent Computer Interface (ACI) to empower SE agents in creating & editing code files, navigating repositories, and executing programs. Additionally, AutoCodeRover (Zhang et al., 2024b) extracts abstract syntax trees in programs, iteratively searches for useful information based on requirements, and generates program patches.
@@
Although these works achieved promising performance, their designs, focusing on local code information, failed to grasp the global context and intricate interdependencies among functions and classes. For example, SWE-agent (Yang et al., 2024) maintains a context window within a code file that allows the agent to scroll up and down. AutoCodeRover (Zhang et al., 2024b) searches functions or classes within the whole repository. Typically, the code comprising a full logic chain for a specific functionality is not arranged sequentially within a single file; rather, it is logically scattered across multiple folders and files. It is difficult to retrieve all relevant code files among maybe thousands of files in a repository, especially starting only from the text in user requirements. This paper argues that a comprehensive understanding of the whole repository becomes the most critical path to ASE. This also is the basis for the multi-file editing functions in existing commercial software like Cursor (Cursor, 2024) and TONGYI Lingma (Alibaba Cloud, 2024).
@@
Undoubtedly, it is challenging to utilize the vast information of an entire repository within LLM. Firstly, a GitHub repository may contain thousands of code files, making it impractical to include them all in the context windows of LLM. Even if it could, an LLM would struggle to accurately capture the code relevant to the objective within such an extensive context. Secondly, the intrinsic logic of how the code execution is distinctly different from the sequence of the code text in a file. For instance, the location where a bug triggers an error message and the actual place that requires modification may not be in the same file, yet they are certainly logically connected.
To address these challenges, we propose LingmaAgent, a novel ASE method deployed in TONGYI Lingma (short for Lingma). Lingma is an IDE-based coding assistant recently developed by Alibaba Cloud and available to users worldwide. Inspired by how human software engineers approach project-level issues, LingmaAgent guides LLM-based agents to first gain a comprehensive understanding of the entire software repository. This approach enables agents to grasp the overall structure and dependencies, thereby enhancing their ability to effectively resolve issues within the broader context of the project.
@@
Specifically, we construct a repository knowledge graph using a top-down approach, organizing the repository into a hierarchical structure tree that provides a clear understanding of code context and scope. This structure is further enhanced by expanding it into a reference graph, capturing intricate function call relationships and facilitating comprehensive dependency and interaction analysis. Subsequently, we propose a Monte Carlo Tree Search (MCTS) based repository exploration method. Specifically, the agents first collect the critical information regarding to the SE task on the repository knowledge graph by the explore-and-exploit strategy. Then, by simulating multiple trajectories and evaluating their reward score, our method iteratively narrows down the search space and guide the agents to focus on the most relevant areas. In addition, to better utilize the repository-level knowledge, we guide the agents to summarize, analyze, and plan for the repository information. Finally, the agents are instructed to manipulate the search API tools to dynamically acquire local information, and fix the real-world issues by generating patches.
We demonstrate the superiority and effectiveness of LingmaAgent through extensive experiments and comprehensive analyses. Using the SWE-bench benchmark (Jimenez et al., 2024), we evaluate our method’s capabilities for issue resolution. Our experiments reveal an 18.5% relative improvement compared to SWE-agent on the SWE-bench Lite benchmark. In production deployment and evaluation at Alibaba Cloud, LingmaAgent automatically resolved 16.9% of in-house issues faced by development engineers and solved 43.3% of problems after manual intervention.
The main contributions of this paper are summarized as follows.
@@
  • We highlight the whole repository understanding as the crucial path to ASE and propose a novel agent-based method named LingmaAgent to solve the challenges.
@@
  • Extensive experiments and analyses demonstrate the superiority and effectiveness of LingmaAgent.
@@ 2.1. LLM-based Software Engineering Agents
In recent years, Large Language Model (LLM) based AI agents have advanced the development of automatic software engineering. AI agents improve the capabilities of project-level software engineering (SE) tasks through running environment awareness (Hong et al., 2023; Wang et al., 2024a; Kong et al., 2024), planning & reasoning (Wang et al., 2024a; Cognition, 2023; OpenDevin, 2023; Luo et al., 2024), and tool construction (Zhang et al., 2024a; Lee et al., 2024). Surprisingly, Devin (Cognition, 2023) is a milestone that explores an end-to-end LLM-based agent system to handle complex SE tasks. Concretely, it first plans the requirements of users, then adopts the editor, terminal and search engine tools to make independent decisions and reasoning, and finally generates codes to satisfy the needs of users in an end-to-end manner. Its promising designs and performance swiftly ignited unprecedented attention from the AI community and SE community to Automatic Software Engineering (ASE) (Yang et al., 2024; Zhang et al., 2024b). For example, SWE-agent (Yang et al., 2024) carefully designs an Agent Computer Interface (ACI) to empower the SE agents capabilities of creating & editing code files, navigating repositories, and executing programs. Besides, AutoCodeRover (Zhang et al., 2024b) extracts the abstract syntax trees in programs, then iteratively searches the useful information according to requirements, and eventually generates program patches. Their designs mainly focus on the local information in the repository, e.g., code files, classes, or functions themselves. Although achieving promising performance, from the insights of the human SE developers, the excellent understanding of the whole repository is a critical path to ASE.
@@ 2.2. Evaluation of LLM-based Software Engineering Agents
Benefiting from the strong general capability of LLMs, LLM-based software engineering agents can handle many important SE tasks, e.g., repository navigation (Zhang et al., 2024a; Wang et al., 2024a), code generation (Hong et al., 2023; Ding et al., 2024a; Ishibashi and Nishimura, 2024; Tang et al., 2024; Rasheed et al., 2024), debugging (Hong et al., 2023; Yang et al., 2024; Zhang et al., 2024b), program repair (Qin et al., 2024; Zhang et al., 2024b; Yang et al., 2024). The existing methods usually regard code generation as a core ability and mainly conduct evaluations on it. Precisely, the code generation test set (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024; Zheng et al., 2023b; Lu et al., 2021) consists of the short problem description, the solution, and the corresponding unit test data. However, with the fast development of LLMs and agents, these datasets are no longer able to comprehensively evaluate their capabilities in the real-world SE tasks. To this end, the repository-level code completion and generation tasks (Liu et al., 2023; Ding et al., 2024b; Du et al., 2024) are presented to evaluate the repository understanding and generation capabilities of LLMs and agents. More recently, SWE team(Jimenez et al., 2024; Yang et al., 2024) develop a unified dataset named SWE-bench to evaluate the capability of the agent system to solve real-world GitHub issues automatically. Specifically, it collects the task instances from real-world GitHub issues from twelve repositories. Consistent with previous evaluation methods, SWE-bench is based on the automatic execution of the unit tests. Differently, the presented test set is challenging and requires the agents to have multiple capabilities, including repository navigation, fault locating, debugging, code generation and program repairing. Besides, SWE-bench Lite (Carlos E. Jimenez, John Yang, Jiayi Geng, 2024) is a subset of SWE-bench, and it has a similar diversity and distribution of repositories as the original version. Due to the smaller test cost and more detailed filtering, SWE-bench Lite is officially recommended as the benchmark of LLM-based SE agents. Therefore, consistent with previous methods (Yang et al., 2024; Zhang et al., 2024b; OpenDevin, 2023), we report our performance on SWE-bench Lite.
@@ 3.1. Overview
We first describe the overall operating process of LingmaAgent, and introduce the stages in detail in the subsequent parts of this section. Given a workspace, LingmaAgent can automatically solve real-world issues. Among them, LingmaAgent involves three key steps, repository knowledge graph construction stage, MCTS-enhanced repository understanding stage, information utilization & patch generation stage. The overall workflow is shown in Figure 1.
@@
Figure 1. The overview of our proposed LingmaAgent. Firstly, we construct the repository knowledge graph is constructed to efficiently represent the code and the dependency in the repository. Subsequently, we empower the agents with the ability of repository understanding by designing the Monte Carlo tree search based repository explore strategy. In addition, we guide the agents to summarize, analyze, and plan to better utilize the repository-level knowledge. Then, they can manipulate the tools to dynamically acquire issue-relevant code information and generate the patches to solve the real-world GitHub issues.
@@
In Repository Knowledge Graph Construction phase, LingmaAgent first builds a repository knowledge graph to represent the entire repository and describe the relationships between entities. This is achieved by parsing the software structure and analyzing it in a top-down manner. The repository is first organized into a hierarchical tree that allows a clear understanding of the context and scope of the code. To facilitate comprehensive dependency and interaction analysis, the tree structure is further extended into a reference graph that captures function call relationships.
@@
Due to the large scale and information complexity of the repository knowledge graph, during the MCTS-Enhanced Repository Understanding phase, LingmaAgent uses the Monte Carlo tree search algorithm to dynamically explore the entire graph. This method focus on discovering key information (i.e., repository functionality and dependency structure) that has a significant impact on issue solving. Through correlation expansion and reference expansion, MCTS simulates multiple trajectories and evaluates their importance, dynamically narrowing the search space and allocating computing resources to the most relevant regions. This targeted navigation enables the model to efficiently access and process important information in the repository, thereby facilitating precise fault localization and informed patch generation.
@@
Inspired by the actual development experience of human programmers, it is necessary to have certain global prior knowledge of the repository before solving specific tasks. Therefore, in Information Utilization & Patch Generation phase, LingmaAgent first summarizes the important information collected in the repository understanding phase to form an experience of the entire repository. Then, in order for LingmaAgent to use the global experience to obtain dynamic information during the problem solving process, we follow AutoCodeRover(Zhang et al., 2024b)’s information retrieval method and use API search tools to extract information in the repository knowledge graph. This includes specific classes and functions and code snippets, etc., to maintain local dynamic knowledge during the task. After collecting enough context, LingmaAgent uses global experience to summarize the currently acquired information to locate faults, generate modified code and return patches that try to resolve the issue.